Consider a portfolio that contains two stocks. Stock "A" has an expected return of 10% and a standard deviation of 20%. Stock "B" has an expected return of -10% and a standard deviation of 25%. The proportion of your wealth invested in stock "A" is 60%. The correlation between the two stocks is 0.

What is the expected return of the portfolio? Enter your answer as a percentage. Do not include the percentage sign in your answer.

Enter your response below rounded to 2 DECIMAL PLACES.

What is the standard deviation of the portfolio? Enter your answer as a percentage. Do not include the percentage sign in your answer.

Enter your response below rounded to 2 DECIMAL PLACES.
Answer #1

Given : weight of stock A (Wa) = 60%

weight of Stock B (Wb) = 1 - 0.60 = 40%

Return from Stock A (Ra)= 10%

Return from stock B (Rb) = - 10%

Standard deviation of A = 20%

Standard deviation of B = 25%

Correlation = 0

Expected return of portfolio = Wa × Ra + Wb × Rb

= 0.60 × 10 + 0.40 × (-10)

= 6 - 4

   = 2

Standard deviation of portfolio

= √ (Wa)^2 (std deviation of A)^2 + (Wb)^2 (std ddeviation of B)^2

= √ (0.6)^2 (0.2)^2 + (0.4)^2 (0.25)^2

= √ 0.36 × 0.04 + 0.16 × 0.0625

= √ 0.0144 + 0.01

= √ 0.0244

= 0.1562 or 15.62